# arXiv:2603.01180v1[astro-ph.HE] 01 Mar 2026

### 01 Mar 2026 v1

@@ 3.3 Radiative properties
Figure 8: shows the bolometric bremsstrahlung luminosity as a function of time for all accretion cases A1–A5. The luminosity is computed for a black hole mass of $M=6.5\times 10^{9}\,M_{\odot}$ and an accretion rate of $\dot{m}=0.01$ (in units of the Eddington rate), corresponding to a physical mass accretion rate of $\dot{M}\approx 9.1\times 10^{24}\,\mathrm{g\,s^{-1}}$ ($\approx 0.14\,M_{\odot}\,\mathrm{yr^{-1}}$). The resulting bolometric luminosities are of the order of $10^{42}$ for case A5   $\mathrm{erg\,s^{-1}}$, yielding a radiative efficiency of $L_{\mathrm{bol}}/(\dot{M}c^{2})\approx 1.2\times 10^{-4}$.
@@
Fig.  8 shows the temporal evolution of the bolometric bremsstrahlung luminosity (erg/sec)for the five cases marked in the legend. Time is measured in years. The bremsstrahlung emission is computed as the volume-integrated emissivity of the plasma using Eq. 25 and primarily contributed by the density and temperature distribution inside the CENBOL region. For the lowest angular momentum case A1 ($l=1.50$), the emission remains very low and nearly constant throughout the simulation. This is consistent with the absence of strong compression and variability inside the CENBOL region as found in Fig.  3 and Fig.  4. For case A2 ($l=1.60$), the luminosity rises slightly and shows small amplitude variability compared to A1. For higher $l$ cases (runs A3, A4, and A5), we observe the bolometric luminosity to increase with $l$. At the same time, all these cases show significant variability; the amplitudes vary by a few factors. Average size of the CENBOL increases with increasing $l$. Thus, increasingly larger volumes with denser matter enhance the bremsstrahlung emission. The variability is caused by the dynamical CENBOL. Fig. 4 already indicate the variability of CENBOL, and observed variability in luminosity is consistent with this.
@@
In passing, we would like to mention that even though our disk-jet systems produce photons of energy $\sim$ MeV, we do not consider the electron-positron pair production. The possibility pair production in high-energy astrophysical sources is not determined solely by photon energy. Instead, it depends critically on the radiation density, which is quantified by a dimensionless parameter (Guilbert et al. 1983):
@@
where $L$ is the luminosity measured in the unit of Eddington luminosity, $D$ is the characteristic size (measured in $r_{g}$ unit) of the emitting region, $\sigma_{T}$ is the Thomson cross-section, and $m_{e}$ is the electron mass. For our case, The Eddington luminosity is
$L_{\rm Edd}=1.26\times 10^{38}\left(\frac{M_{bh}}{M_{\odot}}\right)\simeq 8.2\times 10^{47}\ \mathrm{erg\ s^{-1}}.$
The maximum bolometric luminosity we observe from the result
$L_{\rm bol}\sim 10^{42}\ \mathrm{erg\ s^{-1}}.$
Hence,
$\frac{L_{\rm bol}}{L_{\rm Edd}}\sim 10^{-6}.$
If we consider
$D\sim 2650r_{g},$
we find
$\ell\sim 10^{-9}.$
Since $\ell\ll 1$ (Svensson 1984), the photon–photon optical depth is extremely small. Although the bremsstrahlung spectrum extends up to $\sim 100$ MeV, the radiation field is too dilute for significant pair production. High-energy photons escape without generating a pair cascade.
@@ 4 Summary and Conclusions
In this paper, we study unmagnetized accretion disk-outflow system around a super-massive black hole. These simulations results are equally applicable for stellar mass black holes also provided the radiative efficiency is low. We simulate the accretion of nearly free-falling sub-Keplerian matter which develops a centrifugal barrier due to non-zero angular momentum. A part of the infalling matter is compressed due to this barrier and leaves the system as outflow. In this work, we investigate this mechanism and quantify the outflow by varying the angular momentum of the infalling matter. We use radiative-hydrodynamics simulations for our study, where radiative cooling due to bremsstrahlung emission is included. Our main findings and conclusions are as follows:
@@ Acknowledgements.
We acknowledge the usage of the Kepler Computing facility, maintained by the Department of Physical Sciences, IISER Kolkata. We also acknowledge the usage of IUCAA’s Pegasus Computing facility for conducting a few simulations. Authors also acknowledges the help of Mr. Bungkiu Kissinquinker during the initial phase of this work.

### 13 Apr 2026 v2

@@ 3.3 Radiative properties
Figure 8: shows the bolometric bremsstrahlung luminosity as a function of time for all accretion cases A1–A5. All our simulations are done for a black hole of mass of $M=6.5\times 10^{9}\,M_{\odot}$ with a mass accretion rate of $\dot{m}=0.01$ (in units of the mass Eddington rate), corresponding to a physical mass accretion rate of $\dot{M}\approx 9.36\times 10^{24}\,\mathrm{g\,s^{-1}}$ ($\approx 0.15\,M_{\odot}\,\mathrm{yr^{-1}}$). The resulting bolometric luminosity is of the order of $10^{42}$ for the case A5   $\mathrm{erg\,s^{-1}}$, yielding a radiative efficiency of $L_{\mathrm{bol}}/(\dot{M}c^{2})\approx 4.29\times 10^{-4}$.
@@
Fig.  8 shows the temporal evolution of the bolometric bremsstrahlung luminosity (erg/sec)for the five cases marked in the legend. Time is measured in years. The bremsstrahlung emission is computed as the volume-integrated emissivity of the plasma using Eq. 25 and primarily contributed by the density and temperature distribution inside the CENBOL region. For the lowest angular momentum case A1 ($l=1.50$), the emission remains very low and nearly constant throughout the simulation. This is consistent with the absence of strong compression and variability inside the CENBOL region as found in Fig.  3 and Fig.  4. For case A2 ($l=1.60$), the luminosity rises slightly and shows small amplitude variability compared to A1. For higher $l$ cases (runs A3, A4, and A5), we observe the bolometric luminosity to increase with $l$. At the same time, all these cases show significant variability; the amplitudes vary by a few factors. Average size of the CENBOL increases with increasing $l$. Thus, increasingly larger volumes with denser matter enhance the bremsstrahlung emission. The variability is caused by the dynamical CENBOL. Fig. 4 already indicates the variability of CENBOL, and observed variability in luminosity is consistent with this.
@@
In passing, we would like to mention that even though our disk-jet systems produce photons of energy $\sim$ MeV, we do not consider the electron-positron pair production. The possibility of pair production in high-energy astrophysical sources is not determined solely by photon energy. Instead, it depends critically on the radiation density, which is quantified by a dimensionless parameter (Guilbert et al. 1983):
@@
where $L$ is the luminosity measured in units of Eddington luminosity, $D$ is the characteristic size (measured in $r_{g}$ unit) of the emitting region, $\sigma_{T}$ is the Thomson cross-section, and $m_{e}$ is the electron mass. For our case, the Eddington luminosity is
$L_{\rm Edd}=1.3\times 10^{38}\left(\frac{M_{bh}}{M_{\odot}}\right)\simeq 8.45\times 10^{47}\ \mathrm{erg\ s^{-1}}.$ (29)
The maximum bolometric luminosity we observe from the result $L_{\rm bol}\sim 10^{42}\ \mathrm{erg\ s^{-1}}$. Hence, $\frac{L_{\rm bol}}{L_{\rm Edd}}\sim 10^{-6}$, if we consider $D\sim 2651r_{g}$, we find $\ell\sim 10^{-9}.$ Since $\ell\ll 1$ (Svensson 1984), the photon–photon optical depth is extremely small. Although the bremsstrahlung spectrum extends up to $\sim 100$ MeV, the radiation field is too dilute for significant pair production. High-energy photons escape without generating a pair cascade.
@@ 4 Summary and Conclusions
In this paper, we study unmagnetized accretion disk-outflow system around a super-massive black hole. These simulations results are equally applicable for stellar-mass black holes, also provided the radiative efficiency is low. We simulate the accretion of nearly free-falling sub-Keplerian matter which develops a centrifugal barrier due to non-zero angular momentum. A part of the infalling matter is compressed due to this barrier and leaves the system as outflow. In this work, we investigate this mechanism and quantify the outflow by varying the angular momentum of the infalling matter. We use radiative-hydrodynamics simulations for our study, where radiative cooling due to bremsstrahlung emission is included. Our main findings and conclusions are as follows:
@@ Acknowledgements.
We acknowledge the usage of the Kepler Computing facility, maintained by the Department of Physical Sciences, IISER Kolkata. We also acknowledge the usage of IUCAA’s Pegasus Computing facility for conducting a few simulations. Authors also acknowledge the help of Mr. Bungkiu Kissinquinker during the initial phase of this work.